OpenStudy (anonymous):

Lat \(a_{n}\) be the nth term of a geometric sequence. If a7=32 and a9=8, which of the following must be true? A. a1>0 B. a1-a2>0 C. a2+a3+a4+...+a100 > 0 @RadEn @ganeshie8

OpenStudy (raden):

a_n = a1r^(n-1) a7 = 32 ----> a1r^6 = 32 ... (1) a9 = 8 ----> a1r^8 = 8 ... (2) a1r^8/a1r^6 = 8/32 r^2 = 1/4 r = +-1/2 (r = 1/2 or r = -1/2) now, can you find the value of a1 ?

OpenStudy (raden):

@kryton1212

hartnn (hartnn):

logic : when we divide positive numbers, we get a positive ratio, so, r>0 hence, every term of the series will be positive, so A and C must be true. Math : shown by Redn for B, since a9<a7, all higher terms will be less than lower terms, like \(a_{n+1}<a_n\) so, a1>a2

OpenStudy (anonymous):

wait a moment, i am thinking what did @RadEn has said

OpenStudy (raden):

if we subtitute the possible value of r ( + and -) to one of equation above (1 or 2), i sure there are 2 possible values of a1, are a1>0 or a1< 0 @hartnn

OpenStudy (raden):

so, i think A not always be true

OpenStudy (anonymous):

clear. Why B is true then?

OpenStudy (raden):

i didnt say that --"

OpenStudy (anonymous):

but i think B is true. HAHAHA

OpenStudy (raden):

looks if r > 0 then a1 a>0 too while if r < 0 then a1 < 0 too

OpenStudy (anonymous):

what does it mean?

OpenStudy (raden):

wait, looks hartnn is right :) a1r^6 = 32 if we subtitute r=1/2 or r=-1/2, r^6 must be possitive, thus a1 * (positive number) = 32 a1 must be positive too (in other words a1 > 0) so, A is correct

OpenStudy (anonymous):

lol ..

OpenStudy (raden):

haha, sorry

OpenStudy (anonymous):

how about B ?

OpenStudy (raden):

hmmm... a1r^6 = 32 a1 (1/2)^6 = 32 a1 = 32/(1/2)^6 = 32 * 2^6 = 2048 now, a2 = a1r if we take r = 1/2, obviously a2 be (+), is a2 = a1r = 2048 * 1/2 = 1024 if r = -1/2 then a2 be negative : a2 = a1r = 2048 * (-1/2) = -1024 the 1st possible of a1 - a2 are a1 - a2 = 2048 - 1024 = 1024 (it is positive) the 2nd possible : a1-a2 = 2048 - (-1024) = 3070 (positive too) so, B is correct too :)

OpenStudy (raden):

and like hartnn said above that a_(n+1) < a_n, it means the sum of a2+a3+a4+...+a100 > 0 be true so, A,B, and C all correct

hartnn (hartnn):

Looks like logic is hard to understand initially...

OpenStudy (raden):

but looks we were make mistake again, @hartnn :) for r = 1/2. obviously all terms a1,a2,a3,a4,... and so on be positive and obviously the sum of a2+a3+a4+...+a100 > 0 but if r = -1/2, we get all terms respectively : a1 = 2048 a2 = -1024 a3 = 512 a4 = -128 a5 = 64 a6 = -32 .... so on the option C told us the sum started from a2 (without term a1), so its sum : a2 + a3 + a4 + a5 + a6 + .... + a100 = -1024 + 523 - 128 + 64 - 32 + ... - ? i sure the sum above become negative value (< 0) not positive value (>0) sorry again, @kryton1212 . im not careful this season so, A and B are correct but not always for C